OpenStudy (anonymous):

can someone show me how to factor: -x^4+10x^2-9

OpenStudy (anonymous):

\[x^2(-x^2+10)-9\]

OpenStudy (anonymous):

can you show me the steps? If i wanted the answer I could just use wolfram alfa.

OpenStudy (anonymous):

sorry, i just solved this, and the answer came right to me.

OpenStudy (anonymous):

Thanks for telling me about wolfram alfa. lol

OpenStudy (anonymous):

also, I want to know the factor by grouping method, which I think comes out a little differently.

OpenStudy (anonymous):

oh! You know what, my comp is super slow for some reason, I just got your messages.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but really, can you help with factor by grouping? I keep forgetting...

OpenStudy (anonymous):

sorry. i havent done factoring in months.

OpenStudy (anonymous):

yeah, me neither. Im finding real zeroes in this hmwk assignment and factoring is going a little silly in my head, lol

OpenStudy (anonymous):

Agrreed

OpenStudy (anonymous):

take x^2 as y, -y^2+10y-9=0 multiply the whole thing with -1 y^2-10y+9=0 (y-1)(y-9)=0 the put back x^2 in place of y (x^2-1)(x^2-9)=0 (x+1)(x-1)(x-3)(x+3)=0 I think this is the way, I haven't done this for about 9 years